Identify the quadratic equations as pure, incomplete, or complete
3x² = 42

Answer by Edwin McCravy(20054)

3x² = 42

Is a PURE quadratic equation because it has no term in x.

A COMPLETE quadratic equation 

ax²+bx+c = 0 

has three terms:

(1)  A term in x², "ax²", where a is some number other than 0.
(2)  A term in x, "+bx", where b is some number other than 0.
(3)  A term which is just a number, "+c"

If (1) is missing, it is not a quadratic equation at all, but a linear equation.
If (2) is missing, it is a PURE quadratic equation.
If (3) is missing, it is an INCOMPLETE quadratic equation.  

Example of a COMPLETE quadratic equation: 2x²-5x-7 = 0
Example of an INCOMPLETE quadratic equation:  3x²-4x = 0

Edwin